# Online Graph Learning via Time-Vertex Adaptive Filters: From Theory to Cardiac Fibrillation

### v1

@@ V-C Convergence Across Graph Topologies
V-D Identification of the GSO Topology
Fig. 2 compares the true and estimated GSO matrices across the R, ER, KR, and SBM topologies. For each topology, we show the true GSO $\mathbf{W}$ alongside the mean estimate over 20 realisations, $\text{Mean}(\mathbf{\hat{W}})$, from the best-performing AdaCGP variant (based on lowest steady-state $\text{NMSE}(\mathbf{x}_{h})$). While there is some bias in the estimated element values, particularly in the KR topology, the relative edge weights and overall structure were well preserved.
A comparison of the most common non-zero elements over realisations, $\text{Mode}(\mathbf{\hat{W}}\neq 0)$, to the true elements in $\mathbf{W}$ reveals no false positives as the most frequent outcome, with minimal false negatives. This demonstrates the effectiveness of AdaCGP’s sparse algorithm in explicitly zeroing out GSO matrix elements while identifying most of the true edges.
@@ V-E The Effect of Sparsity Regularisation on Performance
To investigate how estimated network sparsity relates to $\text{NMSE}(\mathbf{x}_{h})$, a practical metric for hyperparameter tuning that does not require knowledge of the true network structure, we experimented with varying the sparsity regularisation strength. Using the best-performing hyperparameters from the AdaCGP variants from before, we randomly sampled elements of $\boldsymbol{\mu}$ uniformly from (0.001, 1] across 5,000 trials per model variant and different graph types with $N=50$ nodes.
@@
VI Assessing the Structure and Stability of Cardiac Fibrillation
While existing applications of Granger causality to fibrillation dynamics [40] use batch processing to capture average propagation patterns, our online algorithm tracks temporal changes in GSO weights and causal structure. This enables assessment of propagation patterns and their stability at both global and local levels, having the potential to inform clinical evaluations of the electrophenotype, where propagation appears chaotic yet maintains quantifiable structure [41, 42, 43]. Such dynamic assessment could identify the stable conduction patterns maintaining arrhythmias, directly supporting diagnosis and treatment strategies.
@@ VI Assessing the Structure and Stability of Cardiac Fibrillation
Figure 4: Global analysis of fibrillation dynamics across varying rotigaptide concentrations. (a) Prediction error $\text{NMSE}(\mathbf{x}_{\Psi})$ over time, showing lower steady-state errors with increasing concentrations. (b) Distribution of Frobenius norms between $\mathbf{W}(t)$ and $\mathbf{W}(t+0.1\text{s})$ in steady-state, indicating more stable dynamics at higher concentrations. (c) Mean Frobenius norm as a function of time lag, demonstrating more gradual changes in graph structure for higher concentrations.
Figure 5: Local stability analysis of fibrillation dynamics. Net edge directions at steady-state start ($t=t^{*}$) and end ($t=T^{*}$) for VF (0 $\mu$M, top) and rotigaptide 50 $\mu$M (bottom). Colour and arrow length denote edge weights. Time series show Out-In Degree for circled region during steady-state, demonstrating alternating source-sink behaviour in VF versus stable source characteristics at 50 $\mu$M.
To demonstrate these capabilities, we analysed optical mapping data of VF from an ex vivo Sprague-Dawley (SD, Charles River, Harlow, UK) rat heart during sequential administration of rotigaptide at concentrations in micromolar ($\mu$M), an anti-arrhythmic drug shown to induce more organised dynamics [43]. All procedures were performed in accordance with the UK Animals (Scientific Procedures) Act 1986 and ARRIVE guidelines, and were approved by the Imperial College London Ethical Review Board (project licences PEE7C76CD and PCA5EE967). Animal procedures conformed to guidelines from Directive 2010/63/EU of the European Parliament on the protection of animals used for scientific purposes. Rats were anaesthetised with 5% isoflurane (95% oxygen mix) and euthanised via cervical dislocation prior to heart explantation.
The experiment records fluorescence signals proportional to cardiac action potential using a 128$\times$80 pixel camera at 1000 Hz. Recordings span 10s during administration of rotigaptide in concentrations of 0, 10, 30, 50, and 80 $\mu$M, where zero represents baseline VF. After following the preprocessing steps in [43], we downsampled the spatial resolution by a factor 4 to $\approx$350 nodes and extracted Hilbert phase angles as graph signals $\mathbf{x}_{t}$, which capture each region’s position in its oscillatory cycle.
AdaCGP (P1 w/ alt debias) is deployed with gradients constrained to adjacent pixels to reflect the local connectivity of cardiac tissue. The model uses parameters $P=3$, $\gamma=1$, $\eta=0.01$, elements of $\boldsymbol{\mu}$ as 0.1, and $\lambda=0.5$, with low forgetting factor enabling rapid adaptation. Stepsizes were calculated using the Armijo rule for automatic selection.
@@
Fig. 4 reveals three aspects of global stability. First, $\text{NMSE}(\mathbf{x}_{\Psi})$ showed lower steady-state errors with increasing rotigaptide concentrations, suggesting more organised and predictable dynamics consistent with the drug’s anti-arrhythmic nature [44]. Second, Frobenius norms between successive GSOs ($\hat{\mathbf{W}}(t)$ and $\hat{\mathbf{W}}(t+0.1\text{s})$) exhibited lower median values and less variability at higher concentrations, indicating more stable global dynamics. Finally, mean Frobenius norm error increased more gradually with time lag at higher concentrations compared to VF (0 $\mu$M) and low concentrations, further supporting enhanced global stability with increased rotigaptide concentration.
Fig. 5 demonstrates local stability through edge patterns at the start ($t^{*}$) and end ($T^{*}$) of steady-state. During VF (0 $\mu$M), the magnitude and direction of net edges varied considerably between $t^{*}$ and $T^{*}$, indicating unstable propagation patterns. The highlighted region’s Out-In Degree ($\sum_{j}\hat{W}_{ij}-\sum_{i}\hat{W}_{ij}$) fluctuated significantly between positive (source) and negative (sink) values, further demonstrating this instability. In contrast, at 50 $\mu$M, edge magnitudes and directions across time points remained largely consistent, suggesting more stable propagation patterns. This stability is quantified by the highlighted region’s behaviour, which maintains consistent source characteristics with minimal Out-In Degree variations that are significantly lower than during VF.
@@ VII Conclusion
We have developed AdaCGP, an online time-vertex adaptive filtering algorithm for tracking time-varying causal graph structures in multivariate time series. Our approach introduced a variable splitting approach for sparse adaptive filtering to efficiently identify the GSO in real-time and reliably estimate its causal elements from streaming data. Experiments across various graph topologies have demonstrated AdaCGP’s superior performance over existing adaptive VAR models in both prediction accuracy and graph structure recovery. Path 1 variants have showed particularly strong performance in causal discovery while maintaining low false positive rates, with prediction error minimisation proving effective for hyperparameter tuning and accurate sparsity estimation. Application to rat VF data revealed AdaCGP’s potential for analysing complex cardiac arrhythmia. The algorithm has successfully tracked changes in cardiac organisation under varying rotigaptide concentrations, providing insights into the structure and stability of fibrillation at both global and local levels. Upon further trials in humans, this approach could help identify critical mechanisms sustaining the disorder, paving the way for personalised treatment strategies and potentially improved outcomes. Future developments could extend our framework to non-linear adaptive filtering and incorporate domain-specific GSO constraints through our variable splitting approach. Such extensions could improve the modelling of non-linear dynamical systems, such as cardiac fibrillation, where non-linear filters could better capture the chaotic dynamics and constrained GSOs, like the anisotropic Laplacian or energy-preserving shift operators [45], would more closely align with established physical models of cardiac conduction [46]. Overall, this work provides a principled and efficient framework for tracking evolving causal relationships in streaming data, bridging the gap between GSP theory and adaptive signal processing for real-time graph structure learning.

### v2

@@ V-D Identification of the GSO Topology
A comparison of the most common non-zero elements over realisations, $\text{Mode}(\mathbf{\hat{W}}\neq 0)$, to the true elements in $\mathbf{W}$ reveals minimal false positives and false negatives. This demonstrates the effectiveness of AdaCGP’s sparse algorithm in explicitly zeroing out GSO matrix elements while identifying most of the true edges.
@@ V-E The Effect of Sparsity Regularisation on Performance
To investigate how estimated network sparsity relates to $\text{NMSE}(\mathbf{x}_{h})$, a practical metric for hyper-parameter tuning that does not require knowledge of the true network structure, we experimented with varying the sparsity regularisation strength. Using the best-performing hyper-parameters from the AdaCGP variants from before, we randomly sampled elements of $\boldsymbol{\mu}$ uniformly from (0.001, 1] across 5,000 trials per model variant and different graph types with $N=50$ nodes.
Fig. 3 shows the $\text{NMSE}(\mathbf{x}_{h})$ versus non-zero elements in $\hat{\mathbf{W}}$ (left column) and $P_{M}$ versus $P_{FA}$ trade-offs (right column), where the x-axis has been uniformly binned and median results displayed. As expected, increasing sparsity regularisation reduces the estimated number of non-zero elements and increases $P_{M}$. Path 1 variants minimised $\text{NMSE}(\mathbf{x}_{h})$ near the true sparsity level, while Path 2 variants overestimated the number of non-zero elements. Path 1 consistently achieved lower $\text{NMSE}(\mathbf{x}_{h})$ and superior $P_{M}$-$P_{FA}$ trade-offs across all topologies, exhibiting near-perfect performance with very low $P_{M}$ and $P_{FA}$. These results demonstrate that optimising NMSE($\mathbf{x}_{h})$ for Path 1 variants leads to accurate network sparsity estimation and thus improved causal discovery. We note that the results in this section serve a different purpose than those in Section V-C and cannot be directly compared. While our previous analysis used optimised hyper-parameters, this section deliberately varies regularisation strength to explore the relationship between sparsity and performance metrics, with results aggregated across multiple configurations including sub-optimal ones for improved visualisation.
Figure 4: Computational complexity analysis of graph topology learning algorithms. Left panels: Scaling of time (top) and memory usage (bottom) with network size, NN. Right panels: Time (top) and memory usage (bottom) as a function of iterations.
V-F Comparison of Runtime and Resource Utilisation
To assess computational efficiency, we measured CPU time per iteration and peak memory usage for the Random topology across network sizes ($N=10$ to $2500$), averaging results over three independent realisations. All algorithms were implemented in NumPy [53] on an AMD EPYC 7513 CPU, and original source code for baselines was used when available. For all auto-regressive methods, we set $P=3$, and all adaptive algorithms employed a fixed step-size. Fig. 4 illustrates these metrics for all algorithms, with empirical scaling exponents derived by fitting polynomial regression models to the highest 50% of tested $N$ values. We excluded AdaCGP’s debiasing step (Algorithm 2) from these experiments, since it is optional and has the same worst-case complexity as Algorithm 1.
The left panels of Fig. 4 show how the computational complexity of each algorithm scales with $N$. AdaCGP variants demonstrate empirical time complexity of $O(N^{2.91})$ for Path 1 and $O(N^{2.89})$ for Path 2, aligning with our theoretical $O(N^{3})$ complexity in Section IV. Despite cubic scaling, our implementation uses efficient matrix operations, achieving competitive runtimes across all tested networks (N=10N=10-25002500) and outperforming several baselines. Memory usage follows similar patterns, with AdaCGP variants maintaining $O(N^{2.00})$ scaling comparable to adaptive baselines. Table II summarises the scaling properties for all algorithms, with VAR+Granger showing highest time complexity, $O(N^{5.81})$, due to causality testing across all potential edges. For GL-SigRep, we report the theoretical time complexity of $O(N^{2.00})$ from [23], as our implementation was not sufficiently optimised for large-scale benchmarking. Additionally, we observe TIRSO’s time complexity of $O(N^{2.05})$ which is significantly lower than its theoretical scaling of $O(N^{3.00})$, likely because the cubic scaling has not become dominant at the network sizes tested.
TABLE II: Scaling of per iteration time and memory complexity. Asterisk denotes theoretical scaling.
Complexity
Algorithm Time Memory
Batch GLasso $O(N^{2.31})$ $O(N^{1.99})$
GL-SigRep $O(N^{2.00})^{\ast}$ -
VAR $O(N^{2.42})$ $O(N^{1.17})$
VAR + Granger $O(N^{5.81})$ $O(N^{3.96})$
Adaptive SD-SEM $O(N^{2.32})$ $O(N^{2.00})$
TIRSO $O(N^{2.05})$ $O(N^{2.00})$
TISO $O(N^{1.98})$ $O(N^{2.00})$
AdaCGP (P2) $O(N^{2.89})$ $O(N^{2.00})$
AdaCGP (P1) $O(N^{2.91})$ $O(N^{2.00})$
The right panels of Fig. 4 demonstrate the constant per-iteration time and memory complexity of adaptive methods, a crucial property that renders them advantageous for the efficient analysis of streaming signals. In contrast, batch methods such as VAR show increasing computational demands as the training window grows (this observation is less apparent for VAR+Granger due to the significant overall cost). As discussed in Section IV, sparsity could be leveraged to yield further improvements in scalability through an implementation of AdaCGP using optimised sparse matrix libraries.
VI Assessing the Structure and Stability of Cardiac Fibrillation
While existing applications of Granger causality to fibrillation dynamics [11] use batch processing to capture average propagation patterns, our online algorithm tracks temporal changes in GSO weights and topology. This enables assessment of propagation patterns and their stability at both global and local levels, having the potential to inform clinical evaluations of the electrophenotype, where propagation appears chaotic yet maintains quantifiable structure [54, 55, 56].
@@ VI Assessing the Structure and Stability of Cardiac Fibrillation
[…]
AdaCGP (P1 + alt debias) is deployed with gradients constrained to adjacent pixels to reflect the local connectivity of cardiac tissue. The model uses parameters $P=3$, $\gamma=1$, $\eta=0.01$, elements of $\boldsymbol{\mu}$ as 0.1, and $\lambda=0.5$, with low forgetting factor enabling rapid adaptation. AdaCGP (P2 + alt debias) uses identical hyper-parameters, differing only in the algorithmic approach as detailed in Section IV. Step-sizes for both variants were calculated using the Armijo rule for automatic selection. To benchmark our approach against established methods in the cardiac electrophysiology literature, we implemented VAR+Granger as a comparative baseline, which has been widely used for functional connectivity estimation in cardiac fibrillation [11, 58]. Unlike our online approach, this offline method was trained on the entire time series data. Individual VAR models ($P=3$) were trained for each pixel, using only time series from adjacent pixels to reflect local connectivity and reduce computational complexity. Granger causality tests were then performed on the resulting VAR coefficients at each pixel to determine significant causal connections.
[…]
@@
Figure 6: Global analysis of fibrillation dynamics across varying rotigaptide concentrations. (a) Prediction error $\text{NMSE}(\mathbf{x}_{\mathbf{h}})$ over time, showing lower steady-state errors with increasing concentrations. (b) Time-lag analysis of Frobenius norm differences between GSOs, demonstrating more stable graph topology estimates at higher concentrations and faster adaptation of AdaCGP P1. (c) Distribution of Frobenius norms between GSOs lagged by 1.0 second during steady-state, indicating more stable dynamics at higher concentrations. (d) Temporal evolution of differences between AdaCGP and VAR+Granger topologies, showing decreased variation with increasing rotigaptide concentration.
[…]
@@
Fig. 6 reveals key aspects of global stability through comparisons between AdaCGP variants and VAR+Granger. First, $\text{NMSE}(\mathbf{x}_{\mathbf{h}})$ in Fig. 6a showed lower steady-state errors with increasing rotigaptide concentrations for both AdaCGP variants, suggesting more organised dynamics consistent with the drug’s anti-arrhythmic nature [59]. Second, Fig. 6c showed Frobenius norms between successive GSOs exhibiting lower median values and variability at higher concentrations, indicating more stable network topologies. Our time-lag analysis in Fig. 6b revealed important methodological differences: while both variants showed similar error patterns at larger time scales (>>1 second), AdaCGP (P1) maintained finite error at short time lags (<<1 second) while AdaCGP (P2) approached zero, confirming Path 1’s greater sensitivity to topology changes.
[…]
AdaCGP’s dynamic assessment offers significant potential for clinical applications in cardiac arrhythmias. For fibrillatory conditions like AF and VF, our algorithm can identify the stability of conduction patterns maintaining the arrhythmia, directly informing treatment strategies. Specifically, by highlighting critical areas of the myocardium, AdaCGP could guide physicians’ radio-frequency catheter ablation decisions regarding precise anatomical targets for intervention. This capability is particularly valuable in redo ablation cases, where previous ablation procedures have created scar tissue that distorts normal cardiac anatomy and alters myocardial electrical properties [60].
For successful translation to clinical practice, several challenges must be addressed. First, validating results in human subjects will require carefully designed clinical studies with appropriate ground truth comparisons. Second, while initial results show promise, further investigation is needed to systematically characterise the algorithm’s performance under various noise conditions typical of clinical recordings. Third, regulatory approval pathways and developing clinically interpretable outputs will be necessary to facilitate adoption by healthcare providers. Addressing these challenges will advance clinical integration in cardiology, potentially improving arrhythmia management and enhancing ablation outcomes.
Beyond cardiac applications, AdaCGP demonstrates strong potential for analysing other complex dynamic systems. In neuroscience, our approach could track dynamical functional connectivity in the brain, with applications in understanding neurological and psychiatric disorders. By identifying abnormal dynamic and transient connectivity patterns, AdaCGP could improve the characterisation of conditions such as schizophrenia, Alzheimer’s disease, and depression [61, 62]. Additionally, it could provide insights into how brain networks reconfigure during cognitive tasks such as memorisation, learning, and decision-making, offering a powerful tool for cognitive research.
@@ VII Conclusion
We have developed AdaCGP, an online time-vertex adaptive filtering algorithm for tracking time-varying causal graph topologies in multivariate time series. Our approach introduced a variable splitting approach for sparsity-aware adaptive filtering to efficiently identify the GSO in real-time and reliably estimate its causal elements from streaming data. Experiments across various graph topologies have demonstrated AdaCGP’s superior performance over existing adaptive VAR models in both prediction accuracy and graph topology recovery. Path 1 variants have showed particularly strong performance in causal discovery while maintaining low false positive rates, with prediction error minimisation proving effective for hyper-parameter tuning and accurate sparsity estimation. Application to rat VF data revealed AdaCGP’s potential for analysing complex cardiac arrhythmia. The algorithm has successfully tracked changes in cardiac organisation under varying rotigaptide concentrations, providing insights into the structure and stability of fibrillation at both global and local levels. Upon further trials in humans, this approach could help identify critical mechanisms sustaining the disorder, paving the way for personalised treatment strategies and potentially improved outcomes. Future developments could extend our framework to non-linear adaptive filtering and incorporate domain-specific GSO constraints through our variable splitting approach. Such extensions could improve the modelling of non-linear dynamical systems, such as cardiac fibrillation, where non-linear filters could better capture the chaotic dynamics and constrained GSOs, like the anisotropic Laplacian or energy-preserving shift operators [63], would more closely align with established physical models of cardiac conduction [64]. Overall, this work provides a principled and efficient framework for tracking evolving causal relationships in streaming data, bridging the gap between GSP theory and adaptive signal processing for real-time graph topology estimation.